A particle m​oving with sn initial velocity of 5m/s is subjected to a uniform acceleration of 2.5m/s2.find the displacement in the next 4sec

Grade:9

1 Answers

Beenu Mathew
7 years ago
u=5m/s
a=5/2m/s2 
t=4s
applying 2 equation of motion
s=ut + 1/2at2
s= 5*4 +1/2 * 5/2 * (4)2
on solving u will get,
s= 40m
HOPE THIS HELPS!!
PLS APPROVE
THANK YOU
JEHOYESJIRA